# Song Form-aware Full-Song Text-to-Lyrics Generation with Multi-Level Granularity Syllable Count Control

### v1

@@ C.5 Evaluation Metrics
Appendix D Success Rate of Generation Model
As our models have to predict <END_*> token properly, there is an possibility of failure. To be specific, the Front-P model has a 37% chance of success on the first try, with the probability of success increasing to 90% within 10 attempts. The Front-S model performs better, with an 89% chance of success on the first attempt and a 99% chance within 10 tries. The remaining models perform well, achieving an average success rate of 97% on the first attempt and about 99.8% within 10 attempts. Overall, Table LABEL:tab:result_gen presents generation results from 6,678 samples in total.
Appendix E Results: Lyrics Infilling
Table LABEL:tab:result_inf presents the results for the infilling task. We used Back-S and Back-P as baselines, which are the same models trained in Table LABEL:tab:result_gen, referring to them as LM-S and LM-P respectively. These LM models consider only the past context and generate solely the masked lyrics. If tokens are not masked, they are directly provided to the model. Since LMs require semantic embedding as input, we provided the model with embeddings of the masked lyrics, replacing the masked segments with [_]. Meanwhile, We refer to the proposed Infilling Language Model as ILMs, followed by -S and -P denoting whether they are trained from scratch or pre-trained model, respectively. Same as in the generation experiments, only the results of the samples successfully inferred by all models within 10 attempts are included in Table LABEL:tab:result_inf, resulting in 9578 samples. The model with the lowest success rate is LM-P, which has a success rate of 97.7%, whereas ILM models has a 100% success rate.
The pre-trained models generally outperform those trained from scratch in terms of PPL, although models trained from scratch show better results on syllable count-based metrics for both LMs and ILMs. LMs exhibits superior performance on the PPL compared to ILM. Additionally, LM-S excels in syllable count-based metrics relative to those of ILM. This suggests that comprehensive attention to the context of the lyrics does not significantly impact syllable count accuracy. However, in terms of the BERT-S, ILMs outperforms LMs, as they are designed to consider the context of the entire lyrics to estimate the masked tokens, thereby enhancing context awareness and capturing semantic content in the infilling task. Especially, ILM-S performs even better than ILM-P in syllable count metrics and BERT-S, which indicates that the models are better trained on the newly introduced special tokens. This result aligns with the findings of the generation experiments.
E.1 Infilling Model Variants
Table LABEL:tab:abl_inf_var explores the impact of changes in the infilling approach. We conducted the following ablation study in comparison with the ILM-S model.
Firstly, we experimented with a model trained using a uniform <MASK> token for all conditioning, instead of using specific tokens like <INF_*><SYL:> for different granularity levels and syllable count annotations (denoted as same mask in Table LABEL:tab:abl_inf_var). In this case, syllable condition tokens are only given directly in ahead of each infilling step.
The results indicated a significant decline in the performance of syllable count generation at both paragraph and word levels, as well as a reduction in the overall BERT-S and PPL.
Next, we trained a model without assigning song form tokens after the <START> token (denoted as no songform). While this change did not significantly affect syllable count performance, it also resulted in a notable decrease in the BERT-S and PPL, similar to the same mask case.
Lastly, we combined both conditions–using the same mask and omitting song form tokens–in training a model. This configuration led to a deterioration across nearly all metrics, confirming the importance of specific masking and song form annotations in lyrics infilling tasks.
Appendix F Preliminary Experiment Setting
In this experiment, we used both gpt-3.5-turbo and gpt-4-turbo-preview to generate lyrics based on specific syllable constraints.
We conducted simple task than main generation experiment, i.e., just generate verse paragraph, because full song template is so complex so chatgpt almost always failed to generate the lyrics within the form.
Prompt for Generation.
The following system role instruction was used to generate lyrics:
You are a lyricist, tasked with creating a song verse. This template outlines the input text and the lyric’s structure, breaking it down into multiple lines. Each line is given with syllable count constraints.
The user input was provided in this format:
Input text: "A summary of the verse in lyrics from the dataset"
Lines with syllable counts = "syllable count for each lines, (e.g. 10, 12, 8, 10)"
Prompt for Re-Generation. If the model produced an incorrect number of lines, the following prompt was used to ask the model to regenerate the lyrics:
Generated lyrics do not match the line counts in the given condition.
Given line count is:
<expected_line_count>,
but generated line count is:
<generated_line_count>.
Please regenerate the lyrics.
This allowed for refinement of the lyrics through up to five regeneration attempts.

### v2

@@ 3.2 Lyrics Infilling
5 Results
5.1 Preliminary Experiment
Large language models (LLMs) such as ChatGPT [34] have demonstrated impressive text generation abilities but often struggle with accurate syllable counting [35]. To evaluate this, we asked an LLM to generate a single paragraph based on input text that specified the number of lines and the syllable count per line. A trial was deemed a failure if the model did not produce the correct number of lines within five attempts. ChatGPT 3.5 achieved a success rate of approximately 38%, while ChatGPT 4.0 reached about 57%. Table 1 summarizes the results for 688 successful samples, indicating that while LLMs are capable of generating lyrics that are contextually appropriate to the input text, they perform poorly in matching the required syllable count.
5.2 Lyrics Generation
In Table 3.2, we present the performance of our models across various settings. For each training method, we report on models trained from scratch as well as those initialized with large-scale pre-trained GPT-2 weights.
Front models serve as our baseline. In these models, the complete generation plan is provided upfront—before the <LYR_START> token—so that the model generates full lyrics without additional guidance during decoding. Essentially, we concatenate all special tokens into a single prompt to generate the output. Back models provide the generation plan to the model during decoding, as shown in Figure 2(a). Both approach merges these methods: it first presents the complete generation plan before the <LYR_START> token and then generates the lyrics using the Back strategy. The suffixes -S and -P indicate whether the model was trained from scratch or initialized with pre-trained weights, respectively.
Table 3.2 shows that pre-trained models achieve lower PPL than models trained from scratch, with the Front and Both approaches performing best. A high PPL indicates that the model struggles to predict the correct tokens. The reason Front and Both models exhibit lower PPL compared to Back models is likely because the model can preview all conditions before generating lyrics, enabling it to make more informed predictions. In contrast, the Back models receive conditioning at each decoding step, limiting their foresight and resulting in higher PPL because they cannot fully anticipate the complete structure. However, given conditions directly at each decoding step could confer advantages in predicting syllable counts. The Both models attempt to integrate the strengths of both Front and Back models but at the expense of increased sample length. Thus, each conditioning approach inherently involves a trade-off.
On the other hand, models trained from scratch consistently outperform pre-trained models in terms of syllable count accuracy, as measured by SCD and SCErr. This may be because scratch models learn to handle special tokens—like syllable tokens—more flexibly. Notably, the Back-S model excels in both syllable metrics and BERT-S performance, indicating its superior ability to capture the input text’s semantics by generating immediately after the text embedding.
5.3 Lyrics Infilling
Table 3.2 presents the results for the infilling task. For baseline comparisons, we use the same models as in the generation experiments—referred to as LM-S and LM-P—which generate masked segments using only past context (with unmasked tokens provided directly). In contrast, our proposed Infilling Language Models (ILM-S and ILM-P) leverage both past and future context. While pre-trained models generally achieve lower PPL, models trained from scratch consistently yield better syllable accuracy. LM models exhibit lower PPL than ILMs, and notably, LM-S demonstrates superior syllable count accuracy. These results suggest that leveraging comprehensive context does not necessarily improve syllable count accuracy. However, ILM-S attains higher BERT-S scores than its counterparts, indicating that using full context during infilling enhances semantic coherence. Overall, ILM-S outperforms ILM-P in both syllable metrics and BERT-S, suggesting that training from scratch better adapts to our specialized token structure. These findings are consistent with our observations from the lyrics generation experiments.
5.4 Song Form Consistency Evaluation
To verify that our model generates distinct lyrics for different song form types, we measure semantic similarity (BERT-S) and structural similarity (normalized Levenshtein Distance, NLD) between paragraphs. We compare only distinct paragraphs within the same song (e.g., verse1 vs. verse2, verse1 vs. chorus1) and compute the average values across all songs.
Figure LABEL:fig:songform_consistency displays the results for the Back-S model in a square matrix format. As these comparisons are made between paragraphs within the same song, we observe that the BERT-S generally exceeds 0.7. Notably, the BERT-S scores for comparisons within the same song form (diagonal values) are higher than those for comparisons between different forms. Similarly, lower NLD values for same-form comparisons indicate that fewer edits are needed between similar segments. Although verses exhibit slightly more variability due to their diverse content, they still show greater similarity within the same form compared to different forms. Overall, these results indicate that the model produces lyrics that are consistent within a given song form and distinct across different forms.
5.5 Ablation Study
Table LABEL:tab:abl_inf_var examines the impact of modifications to our infilling strategy compared to the baseline ILM-S model. First, we train a variant using a uniform <MASK> token for all conditioning instead of specialized tokens (e.g., <INF_*><SYL:>) for different granularities; we denote this variant as same mask in Table LABEL:tab:abl_inf_var. In this setup, syllable condition tokens are provided directly at each infilling step. The results show a significant decline in syllable count performance at both the paragraph and word levels, along with reductions in overall BERT-S and increased PPL. Next, we train a model without including song form tokens after the <START> token, denoted as no songform. While this change has little effect on syllable count accuracy, it also results in a noticeable drop in BERT-S and a rise in PPL, similar to the same mask condition. Finally, we combine both modifications—using a uniform mask and omitting song form tokens—in a single model. This configuration degrades performance across nearly all metrics, confirming the importance of our specialized masking strategy and the inclusion of song form annotations for effective lyrics infilling.
6 Conclusion
Our study proposes an effective approach to automated lyrics generation and infilling that leverages multi-level granularity for precise syllable control and effective song form management, all conditioned on an arbitrary input text. This method enhances the flexibility and structural accuracy of the generated lyrics. In future work, we plan to incorporate additional controls—such as genre tags and rhyme schemes—to further improve the adaptability and creative potential of automated lyrics composition.
7 Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No. RS-2024-00461617, 40%], Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [No. RS-2022-II220320, 2022-0-00320, 50%], [No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), 5%], and [No.RS-2021-II212068, Artificial Intelligence Innovation Hub, 5%]